Take 4 coins and toss them a total of 32 times. Each time, record the number of heads observed. For the report, list the numbers from 1 to 32 and record how many heads were observed on each toss. [3 points frequencies as the probability a probability distribution, using the relative
Answer #1

Solution:

   We have to toss 4 coins at a time for 32 times and note down the results for each outcomes and find out how many times heads occurs. Thus we get following results when we tossed 4 coins 32 times.

Sr. No. 1st coin 2nd coin 3rd coin 4th coin No. of Heads.
1 H T T H 2
2 T T H T 1
3 H H T T 2
4 T T T T 0
5 H H T H 3
6 T T T H 1
7 H H T T 2
8 T T T H 1
9 T T H H 2
10 H H H T 3
11 H H H T 3
12 T T H T 1
13 T T H T 1
14 T H T T 1
15 T H T H 2
16 T H H H 3
17 T H H H 3
18 T H T T 1
19 T T H T 1
20 T T T T 0
21 T H T H 2
22 H T H T 2
23 T T H H 2
24 T H H H 3
25 T H H T 2
26 T H H T 2
27 H T T T 1
28 H T T H 2
29 T H H H 3
30 T H H T 2
31 T H T H 2
32 T T H H 2

probability distribution of number of heads:

X= Number of Heads f =frequency P(X)
0 2 2/32=0.06250
1 9 9/32=0.28125
2 14 14/32=0.43750
3 7 7/32=0.21875
4 0 0/32=0.00000
N = 32 \sum P(X)=1.00000